# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 1.4.2 What we adopt and what cannot adopt for black box polynomial root-finding
On the other hand, the root-finder of [76] cannot help solve even Problem 1. Indeed, repeatedly deflating $p$, by peeling off large roots, can accumulate errors that grow too quickly, together with the coefficient length of the factors (notice, e.g., the swell of the coefficients of $p(x)=x^{d}+1$ in transition to its factors of degree $d/2$, say). This is true even if one performs deflation in the black box setting implicitly, based on Eqn. (3.3) of Sec. 3. Moreover, even in that setting the coefficient swell in deflation is translated into swelling of the values of the high order derivatives involved in the algorithm of [76].
Even if the algorithm of [76] worked for Problem 1 it would only solve the well-known and much easier problem of root-finding for a real-rooted polynomial (cf. [93, 25, 33]). Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, which would by a factor of $b$ exceed the upper bound of [96, 103] for general polynomial root-finding.
Recursive combination of [76] with the Inverse Power Iterations [48] could lead to approximation of all eigenvalues of a symmetric (although not unsymmetric) matrix, together with its eigenvectors. One can be challenged to estimate the increased complexity of this extended computations, but we study other approaches.
By applying a root-finder to Given a polynomials $p$ of (1.1) given with its coefficients We accelerate known algorithms for Problem 0 by a factor of $m/\log^{2}(m)$ by incorporating fast multipoint polynomial evaluation algorithms of [42, 84, 68, 81], but we could not use them if we successively approximated zeros of $p$ or matrix eigenvalues by using deflation. Thus, instead of applying high order Newton’s iterations of [76] and deflation, we recall and extend subdivision iterations of [149, 52, 128, 101].
@@ 1.5 Our complexity estimates
$m=m(\mathbb{D},\beta)$ for $\beta>0$ is the number of roots of polynomial $p(x)$ in the $\beta$-neighborhood of a domain $\mathbb{D}$ on the complex plane.
Main Theorems 1-4 below hold for $m$ of Def. 1.1, any fixed constant $\beta>0$, and either $p(x)=\det(xI-M)$ in the case of Problems 2 and 2 or $p(x)$ of (1.1) otherwise.
@@
(Main Theorem 2.) One can solve Problems 2 and 2 at Las Vegas randomized bit operation cost
@@
for $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$, the Frobenius norm $||M||_{F}$ of a matrix $M$, and $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM). [^4] The current records are about 2.7734 for exponents of feasible MM [90, 107, 109, 72, 57], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [6, 37]. The exponents below 2.77 have only been reached at the expense of blowing up the size of MM by immense factors.
@@
(Main Theorem 3.) One can solve Problems 0 and 0 at Las Vegas randomized bit operation cost $\tilde{O}(bdm))$ for a polynomial $p$ of (1.1) normalized such that $||p(x)||_{1}:=\sum_{i=1}^{d}|p_{i}|=1$.
Random Root Model: The zeros of $p$ (in particular for $p$ denoting the characteristic polynomial of a matrix $M$) are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large $R\in\mathbb{R}$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.555(i) The assumption of the model is a little stronger than that of our Def. 24.1, supporting such a decrease in [96, 103]. (ii) Random coefficient models may seem more natural for polynomial root-finding but are useless because no known probability distribution withstands shift of the variable $x$.
@@
(Main Theorem 4) The upper bounds of Thms. 1.1-1.3 on computational precision and bit operation complexity can be decreased by a factor of $d$ under the Random Root Model.
In view of Remarks 1.1 and 25.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
2 Accelerated subdivision polynomial root-finders: overview and reduction to soft exclusion/inclusion tests
For a polynomial $p$ given in monomial basis (with its coefficients) the deterministic algorithm of [96, 103] supports our Main Theorems 3 and 4, but not our Main Theorems 1 and 2. This is because the algorithm of [96, 103] cannot be applied to a black box polynomial, unlike our present Las Vegas randomized algorithms, which we outline next and elaborate upon in Parts II–IV. Namely, in the next two sections we reduce our root-finders to recursive application of exclusion/inclusion (e/i) test, which decides whether a fixed complex disc contains roots or not. In Secs. 3.3.2 and 3.4 we outline our two approaches to devising e/i tests for a black box polynomial and unit disc $\mathcal{D}=D(0,1)$. We cover these two approaches in some detail in Part II, made up of Secs. 612, and Part III, made up of Secs. 1318, respectively, and describe yet another fast e/i test in Sec. 23.4. Our root-finders of Parts II and III incorporate these e/i tests and are quite fast but still only support root approximation (rather than isolation) and the complexity estimates exceeding those of our Main Theorems by a factor of $b$. In Part IV, made up of Secs. 1923 and outlined in Sec. 4, we fix both deficiencies, thus proving our Main Theorems. In Sec. 5.4 we list some tentative directions to further progress.
@@ 2.1 Subdivision root-finders and reduction to e/i tests
We depart from the popular subdivision root-finders, traced back to [149, 52, 128, 101]. They extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction.
Next we outline them.
(A) First let a square on the complex plane contain precisely $m>0$ roots (with multiplicities) and let us approximate them. Call the square suspect, recursively divide every suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test in order to decide whether the square contains a root or does not. If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig. 1a). In Sec. 9 and then in our Part IV we use a natural extension of e/i tests to $m$-tests, which decide whether a disc in the complex plane covering at most $m$ roots does or does not contain $m$ roots.
@@
A subdivision iteration applied to a square with $m$ roots (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every root within the distance of one half of their diameter.
@@
Let subdivision iterations be applied to a square $S$ with side length $\Delta$ and diameter $\Delta\sqrt{2}$ containing precisely $m$ roots. Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta/2^{k-0.5}$ all the $m$ roots lying in $S$, so that $k\leq\lceil\log_{2}(\frac{\Delta}{2\epsilon}\sqrt{2})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
(B) Seeking roots in any fixed domain $\mathbb{D}$ in the complex plane, we apply subdivision iterations to its minimal covering square $S$ and while performing them also discard all squares that do not intersect the domain $\mathbb{D}$, even if they contain roots.
Then the $(\Delta_{k}\sqrt{2})$-neighborhood of $\mathbb{D}$ covers the union $\mathbb{U}_{k}$ of all suspect squares defined after the $k$th subdivision iteration, while $\Delta_{k}=\Delta_{0}/2^{k}$ fast converges to 0 as $k$ increases. Hence the neighborhood fast converges to the domain $\mathbb{D}$: the width of the neighborhood decreases by a factor of $2^{k}$ in $k$ subdivision iterations, still at the price of performing at most $4km$ e/i tests.
@@
[Real and near-real root-finding.] In the case of Problem 0 of Example 1.2, generalize bisection as follows: (i) recursively partition (subdivide) every suspect rectangle into two congruent rectangles, to each apply e/i test, and then discards a rectangle if it contains no roots but call it suspect and subdivides it otherwise, (ii) stop where the rectangles have diameter at most $2\epsilon$, so that their centers approximate the roots within $\epsilon$.
@@ 2.2 Subdivision iterations with soft e/i tests
Instead of testing whether a suspect square $S$ contains some roots, Henrici in [52] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this. Furthermore, in the presence of rounding errors no e/i test can decide whether a root lies in or outside a disc if it actually lies on or very close to its boundary circle. To address both problems we fix $\sigma>1$ and apply soft e/i tests.
@@
[See the end of Sec. 6 for a generalization, covering in particular soft $m$-tests.] Given a real $\sigma>1$, a polynomial $p$, a square $S$, and its minimal covering disc $D(c,\rho)$, the $\sigma$-soft e/i test reports either exclusion or $\sigma$-soft inclusion. Exclusion certifies that the square $S$ contains no roots and therefore should be discarded from subdivision process. $\sigma$-soft inclusion certifies that the disc $D(c,\sigma\rho)$, said to be the $\sigma$-cover of $S$, contains a root; then the square $S$ is said to be suspect and is subdivided into four congruent sub-squares in the next subdivision step. Criteria for both exclusion and $\sigma$-soft inclusion can hold simultaneously (see Fig. 1b), but the test ends as soon as it verifies any of them.
@@
Let $\gamma(\sigma)$ denote the maximal number of $\sigma$-covers at a fixed subdivision step that share a complex point, e.g., a root. Then $\gamma(\sigma)$ is in $O(1)$ if so is $\sigma$, and $\gamma(\sigma)\leq 4$ if $\sigma<\sqrt{2}$.
@@
Cor. 2.1 still holds if e/i tests are replaced with $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$, while a subdivision root-finder using $\sigma$-soft e/i tests involves $O(km)$ such tests if $\sigma$ is a constant.
@@ 3 Fast soft e/i tests
3.1 Background: Newton’s Inverse Ratio and its approximation
We reduce e/i tests of subdivision root-finders to the multipoint evaluation of the ratio
@@ 3.1 Background: Newton’s Inverse Ratio and its approximation
which we call Newton’s inverse ratio. To prove the latter well-known and much used equation, differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$.
The arithmetic operations (hereafter ops) involved in our evaluation of NIR$(x)$ are more numerous than the other ops involved in our root-finders, even in the case where $p(x)=x^{d}$.
By following [76] we fix a complex $\delta$, with $|\delta|$ small in context, and approximate $p^{\prime}(x)$ with the first order divided difference $\frac{p(x)-p(x-\delta)}{\delta}$, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$, by virtue of Taylor-Lagrange’s formula. Hence $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$, and
@@
We deduce the same bound $O(b\log(d))$ of [76] on computational precision of our root-finders, but its derivation is greatly simplified versus [76] because we only use the first order derivatives and the first order divided differences.
Unlike [76] we estimate bit-operation cost of our root-finders and allow random choice of $x$ to control the value of the ratio $\frac{|p(x)-p(x-\delta)|}{p(x)\delta}$ with a high probability (whp). This is Las Vegas randomization, and so is any randomization in polynomial root-finding, as we show next.

### v37

@@ 1.4.2 Progress at FOCS 2016 and its limitations
Besides proposing novel root-finder for a black box polynomial with extension to approximation of an eigenvalue of a matrix, Louis and Vempala in [59] presented non-trivial techniques for bounding precision of root-finding computations. Extension of these techniques helped us simplify dramatically the estimation of the bit operation complexity of our root-finders compared to those of [106, 78, 84, 52], which bypass estimation of computational precision.
On the other hand, direct extension of the root-finder of [59] cannot help solve even Problem 1. Indeed, repeatedly deflating $p$, by peeling off large roots, can accumulate errors that grow too quickly, together with the coefficient length of the factors: notice, e.g., the swell of the coefficients of $p(x)=x^{d}+1$ in transition to its factors of degree $d/2$, say. This is true even if one performs deflation in the black box setting implicitly, based on Eqn. (3.3) of Sec. 3. Indeed, the high order derivatives involved in the algorithm of [59] are proportional to the coefficients and together with them swell in deflation.
Even if deflation succeeded, the algorithm of [59] would have only solved the well-known and much easier problem of root-finding for a real-rooted polynomial (cf. [75, 18, 25]). Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, exceeding by a factor of $b$ the upper bound of [78, 84] for general polynomial root-finding.
Furthermore, we accelerate the solution of Problem 0 by a factor of $m/\log^{2}(m)$ by incorporating fast multipoint polynomial evaluation of [32, 67, 52, 64], but we would lose this chance if we successively approximated the zeros of $p$ or matrix eigenvalues by using deflation.
Thus, instead of applying high order Newton’s iterations of [59] and deflation, we recall and advance subdivision iterations of [120, 39, 100, 82].
@@ 1.5 Our complexity estimates
Unlike the root-finders of [106, 78, 84, 52], seeking all $d$ roots, of [59], seeking a single absolutely largest root and of [78, 84], seeking all $d$ roots or all roots in a disc well isolated from the external roots, we approximate the roots in any domain $\mathbb{D}$ in the complex plane. Root approximation within $\epsilon$ in a domain $\mathbb{D}$ involves the roots in its $\epsilon$-neighborhood; our complexity estimates grows with the number of roots in a little larger but very small neighborhood of $\mathbb{D}$.
@@
For a polynomial $p=p(x)$, a domain $\mathbb{D}$ in the complex plane having a diameter $\Delta$, and a positive $\beta$ such that $\beta\Delta>2\epsilon$, write $m=m(p,\mathbb{D},\beta)$ if precisely $m$ roots of $p(x)$ lie in the $\beta\Delta$-neighborhood of $\mathbb{D}$.
We will prove our Main Theorems 1-4 below for $m$ of Def. 1.1, any fixed constant $\beta>2\epsilon/\Delta$, and either $p(x)=\det(xI-M)$ in the case of Problems 2 and 2 or $p(x)$ of (1.1) otherwise. [^4] Hereafter, whenever we say that subdivision iterations are applied to a input domain $\mathbb{D}$ of diameter $\Delta$ that contains $m$ roots, we mean that so does its $\beta$-neighborhood.
@@
(Main Theorem 2.) One can solve Problems 2 and 2 at a Las Vegas randomized bit operation cost
@@
for $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$, the Frobenius norm $||M||_{F}$ of a matrix $M$, and $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM). [^5] The current records are about 2.7734 for exponents of feasible MM [73, 88, 90, 56, 42], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [122]. The exponents below 2.77 have only been reached at the expense of blowing up the size of MM by immense factors.
@@
(Main Theorem 3.) One can solve Problems 0 and 0 at a Las Vegas randomized bit operation cost $\tilde{O}(bdm))$ for a polynomial $p$ of (1.1) normalized such that $\sum_{i=1}^{d}|p_{i}|=1$.
Random Root Model: The zeros of $p(x)$, for $p(x)=\det(xI-M)$ in Problems 2 and 2 and for $p(x)$ of (1.1) otherwise, are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large real $R$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.666(i) The assumption of the model is a little stronger than that of our Def. 22.1, supporting such a decrease in [78, 84]. (ii) Random coefficient models fails at the basic step of the shift of the variable $x$, which no known probability distribution withstands.
@@
(Main Theorem 4.) Under the Random Root Model one can decrease by a factor of $m$ the upper bounds of Thms. 1.1-1.3 on the computational precision and bit operation complexity.
In view of Remarks 1.1 and 23.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
2 Fast subdivision root-finding: overview and soft e/i tests
We support Main Theorems 1 – 4 with our accelerated subdivision root-finders. They are reduced to recursive application of soft exclusion/inclusion (e/i) tests, which decide whether a fixed complex disc or its small neighborhood contains roots. In Sec. 3 we outline our two approaches to devising e/i tests for a black box polynomial and the unit disc $\mathcal{D}=D(0,1)$. We expand these two outlines in Part II, made up of Secs. 610, and Part III, made up of Secs. 1115, respectively. Our root-finders of Parts II and III only approximate (rather than isolate) the roots and support complexity estimates that are inferior to those of our Main Theorems by a factor of $b$. In Part IV, made up of Secs. 1621 and outlined in Sec. 4, we fix both deficiencies, thus proving the Main Theorems. In Sec. 5.5 we further comment about the contents of our paper, including its Part V, made up of Secs. 2225.
@@ 2.1 Subdivision root-finders and reduction to e/i tests
We depart from popular subdivision root-finders, traced back to [120, 39, 100, 82]. They extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction. Next we outline them.
(A) First consider root-finding in a square on the complex plane. Call that square suspect, recursively divide it and every other suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test to decide whether the square contains a root or does not. If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig. 1a). In Def. 2.1 we generalize e/i tests to $\ell$-tests, which decide whether a disc in the complex plane contains $\ell$ or fewer roots provided that it contains at most $m$ roots. E/i test is a 1-tests, and whenever we need $\ell$-tests for $\ell>1$ (e.g., in Sec. 16 and Part IV), we reduce them to e/i tests (see Sec. 18 and 13.4).
@@
A subdivision iteration applied to a square with $m$ roots in its $\beta$-neighborhood (cf. Def. 1.1 and Remark 2.1) (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every root within the distance of one half of their diameter.
@@
Let subdivision iterations be applied to a square $S$ with side length $\Delta$ and diameter $\Delta^{\prime}:=\Delta\sqrt{2}$ containing precisely $m$ roots. Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta^{\prime}/2^{k+1}$ all the $m$ roots lying in $S$, so that $k+1\leq\lceil\log_{2}(\frac{\Delta^{\prime}}{\epsilon})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
(B) Seeking roots in any fixed complex domain $\mathbb{D}$, we apply subdivision iterations to its minimal covering square $S$ with diameter $\Delta^{\prime}$ and while performing an iteration also discard all squares not intersecting the domain $\mathbb{D}$, even if they contain roots. The following proposition bounds the overall number of e/i tests performed in $k$ iterations.
Proposition 2.1.
Let $k$ subdivision iterations be applied to a complex domain $\mathbb{D}$, let precisely $m_{i}$ roots lie in the $\Delta_{i}^{\prime}$-neighborhood of $\mathbb{D}$ for $\Delta^{\prime}_{i}$ denoting the diameter of a suspect square at the $i$-th subdivision iteration, and let precisely $m_{i}$ e/i tests be performed at the $i$-thiteration, so that
$m_{i}\leq 4^{i+1},~{}i=0,1,\dots,k-1.$ (2.1)
Then
(i) the $(\Delta_{h}^{\prime})$-neighborhood of $\mathbb{D}$ covers the union of all suspect squares processed in the $h$-th subdivision iteration, while $\Delta_{h}^{\prime}=\Delta_{0}^{\prime}/2^{h}$ fast converges to 0 as $h$ increases; in particular, for any fixed $\beta$ of Def. 1.1 we have $\Delta^{\prime}_{h}\leq 0.5\beta\Delta^{\prime}_{0}$ already for $h=O(1)$.
(ii) The first $h$ iterations involve at most $\frac{4^{h+1}}{3}$ e/i tests, that is, $O(1)$ tests if $h=O(1)$.
Proof.
We only prove claim (ii). Eqn. (2.1) implies that $\sum_{i=0}^{h-1}m_{i}\leq\sum_{i=0}^{h-1}4^{i+1}=\frac{4^{h+1}}{3}$. ∎
@@
Every root is defined up to within $\epsilon$, and so at the $h$-th subdivision iteration we decide whether we should discard a square or call it suspect and keep subdividing it depending on the roots that lie in the $(\Delta_{h}^{\prime}+\epsilon)$-neighborhood of $\mathbb{D}$. Def. 1.1 for $\Delta:=\Delta_{0}^{\prime}$ implies that $\Delta^{\prime}_{h}\epsilon\leq\beta\Delta$ already for some $h=O(1)$, and so, the outputs of all e/i tests, except for $h=O(1)$ initial ones, depend only on the roots that lie in the $\beta\Delta$-neighborhood of $\mathbb{D}$.
Remark 2.2.
[Real and near-real root-finding.] In the case of Problem 0 of Example 1.2, generalize bisection root-finding as follows: (i) recursively partition (subdivide) the real line segment defining every suspect rectangle into two halves and accordingly subdivide the rectangle into two congruent rectangles; to each of them apply e/i test, and then discard a rectangle if it contains no roots but call it suspect and subdivide it otherwise, (ii) stop where a rectangle has diameter at most $2\epsilon$, so that its center approximates within $\epsilon$ all roots lying in it.
@@ 2.2 Subdivision iterations with soft e/i tests
Instead of testing whether a suspect square $S$ contains some roots, Henrici in [39] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this. Furthermore, in the presence of rounding errors no e/i test can decide whether a root lies in or outside a disc if it actually lies on or very close to its boundary circle. We address both problems by fixing $\sigma>1$ and applying soft e/i tests, which are a special case for $\ell=1$ of the following more general soft $\ell$-tests. [^7] Besides 1-tests (that is, e/i tests) we only use $\ell$-tests for sub-discs of discs containing at most $\ell$ roots and well-isolated from the external roots.
@@
For $p(x)$ of (1.1), $\sigma>1$, and an integer $\ell$, $1\leq\ell\leq d$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (in particular an e/i test turns into a 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^8] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf. Fig. 1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds. #$\#(D(0,1))<\ell$. An $\ell$-test applied to the polynomial $t(y)$ of (1.4) is said to be an $\ell$-test for the disc $D(c,\rho)$ and/or the circle $C(c,\rho)$ and also to be an $\ell$-testc,ρ.
@@
Let $\gamma(\sigma)$ denote the maximal number of $\sigma$-covers at a fixed subdivision step that share a complex point, e.g., a root. Then $\gamma(\sigma)\leq 4$ if $\sigma<\sqrt{2}$ and $\gamma(\sigma)$ is in $O(1)$ if so is $\sigma$.
@@
Cor. 2.1 and the observation of Remark 2.1 still hold if e/i tests are replaced with $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$, while a subdivision root-finder using $\sigma$-soft e/i tests involves $O(km)$ such tests if $\sigma$ is a constant.
@@ 3 Fast soft e/i tests
3.1 Background: Newton’s Inverse Ratio
We reduce e/i tests of subdivision root-finders to multipoint evaluation of the ratio
@@ 3.1 Background: Newton’s Inverse Ratio
said to be Newton’s inverse ratio. To prove the latter well-known equation, differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$.
The ops involved in our evaluation of NIR$(x)$ are more numerous than the other ops involved in our root-finders, even in the case where $p(x)=x^{d}$.
By following [59] we fix a complex $\delta$, for $|\delta|$ small in context, and approximate $p^{\prime}(x)$ with the first order divided difference $\frac{p(x)-p(x-\delta)}{\delta}$, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$, by virtue of Taylor-Lagrange’s formula. Hence $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$, and